OpenStudy (lucaz):

solution of (x-2)(x+5)<=0 ?

OpenStudy (anonymous):

zeros are at \(-5\) and \(2\) and since \(y=(x-2)(x+5)\) is a parabola that faces up, it is negative between the zeros i.e. if \(-5\leq x\leq 2\)

OpenStudy (lucaz):

right, I forgot the graph. thanks!

OpenStudy (agent0smith):

Yep, a graph helps you to see where the graph is below the x-axis.

OpenStudy (lucaz):

yeh